The quadratic equation whose one root is square root of -47+8√-3 is

Grade:12

1 Answers

Anuj
6 years ago
If it is a square root of given quantity, the given quantity can be broken into square of something.-47+8√(-3)=-47+2×4√(-3)×1=(4√-3)²+(1)²+2×4√(-3)×1={1+4√(-3)}²So one root is_ 1+4√(-3)& The other_. 1-4√(-3).So the equation-x²-2x+49=0